40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 643     Entry time: Mon Jul 7 19:15:38 2008
Author: Alberto 
Type: Update 
Category: General 
Subject: Optics alignement on the ABS length experiment 
Today I started setting up the PLL instruments to lock the frequency of the NPRO beam to the IFO beam. with no need of a new alignment after the weekend I was able to see the beat again, although this time I found at a different temperature of the NPRO laser of about 54 degrees (vs 51 of the last time).
I've got the Marconi as local oscillator (LO), the mixer Koji suggetsed, the SR560 and a 5 MHz low pass filter to cut the 33, 66 and 99 MHz present in the output signal from the PD. The filter worked well and I was able to single out only the beat resonance from the power spectrum.
In the attempt to enhance the amplitude of the beat, as Koji suggested, I tried to work on the alignment of the steering mirrors. While I was doing that, for some reason the pre-modecleaner lost the alignment and I had to ask John to help me lock it again. during the process I lost the old alignment but at the end I got a new one, apparently (from the camera) even better than the other. Although after that the beats were gone. Actually after the lock-in of the PMC the IFO beam didn't look as good as before, so it might be also for that reason.

I'll try again tomorrow, after that probably tonight Rob is going to reset the alignments of the interferometers.
ELOG V3.1.3-